# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v4

@@ 3.2 Model
$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{\frac{1}{p\,\theta(B)}+\frac{1-p}{p\,N_{1}(A)}\left(1+\frac{N_{1}(B)\,N_{1}(A\backslash B)}{\theta(B)^{2}}\right)}$
and the plug-in estimator is
$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)\left(1+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{0}(B)^{2}}\right)}$
See Section 7.3.3 for details.
3.3 Data Analysis
@@
We use the methodology described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City. According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not. Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349. Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
@@
Estimates for select subregions are listed in Section 7.1, Table 1. We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan. Another quarter is located in West and North Queens. We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
@@
We check these estimates by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024). According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 152 street vendors total with a standard error of 54. We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
4. Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is accurate in large samples (i.e., consistent). A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. This assumption is necessary to derive the standard errors.
In this section, we examine each assumption. Since neither can be evaluated from the survey data alone, we consider additional information. In Section 4.1, we compare the spatial distribution of the respondents to data from the New York City Office of Administrative Trials and Hearings and the American Community Survey. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets. [^5] The results of this section can also be derived from a spatial Negative Binomial Lévy process in a manner that mirrors Section 3.2. The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
4.1 Representativeness
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative (i.e., the response probability is spatially uncorrelated with the mean measure). If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem.
If this assumption is violated, then $\hat{\theta}(A)$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}(A)$ would likely underestimate $\theta(A)$ even in large samples.
We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid. We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
Nevertheless, it is possible that canvassing operations systematically missed locations, and these missed locations may be revealed by examining other data sources. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S. Census Bureau (accessed 2022-12-22).
These comparisons are visualized in Section 7.2, Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels), although there are several small discrepancies. For example, the OATH data suggest a larger percentage of vendors work in South Brooklyn (Coney Island and Brighton Beach) than captured by the Street Vendor Project survey, and the ACS data suggest a larger percentage of street vendors live in South Queens (Rockaway). However, both the OATH and ACS data suggest these locations reflect only a small percentage of New York City vendors, and thus any resulting undercount is likely to have little impact on our overall estimate.
We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated. That is because the undercount in some areas may be compensated by an overcount in others (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
4.2 Overdispersion
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because the consistency of $\hat{\theta}(A)$ holds under a wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated. [^6] The ratio estimator may no longer be asymptotically efficient as well. In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies it is asymptotically efficient. However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
@@ 4.2 Overdispersion
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925). That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(B)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\mu_{i}(B)=\exp(\lambda_{i})\,M(B)$ and variance $\mu_{i}(B)\bigr{(}\mu_{i}(B)-M(B)\bigr{)}\,/\,M(B)$. i.e.,
$N_{i}(B)\sim\ \text{Negative Binomial}\left(\mu_{i}(B),\ \mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
where we parameterize the negative binomial distribution by its mean and variance. In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details. [^7] Note that we define the negative binomial and negative hypergeometric distributions as the total number of trials required to get a predetermined number of successes. The negative binomial assumes sampling with replacement, while the negative hypergeometric distribution assumes sampling without replacement.
To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey so that
$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\mu_{i}(B),\ p\,\mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
The conditional distribution of $n_{1}(B)$ given $N_{1}(B)$ is then
$n_{1}(B)\ |\ N_{1}(B)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(B)-1,\ M(B)-1,\ p\,M(B)\bigl{)}$
with conditional mean and variance
$\mathbb{E}\left[n_{1}(B)\ |\ N_{1}(B)\right]=p\,N_{1}(B)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(B)\ |\ N_{1}(B)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(B)$
so that under increasing domain asymptotics, the ratio estimator
$\hat{\theta}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
is asymptotically normal with mean $\theta(A)=\mu_{0}(A)$ and standard error
$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{w_{0}\left(\frac{1}{p\,\theta(A)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$w_{0}=\dfrac{\theta(A)-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.3.4 for details. Note that the weights $w_{0}$ and $w_{1}$ depend on $A$, although this dependence is not reflected in the notation.
The standard error derived in this section for $\hat{\theta}(A)$ is identical to the standard error for $\hat{\theta}(A)$ from Section 3.2 except for the weights $w_{0}$ and $w_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.

### v5

@@ 3.2 Model
$\tilde{\text{SE}}[\tilde{\tau}(B)]=\dfrac{N_{1}(A)\,n_{0}(B)}{n_{1}(A)}\sqrt{\dfrac{1}{n_{0}(B)}+\frac{1}{n_{1}(A)}-\frac{1}{N_{1}(A)}+\dfrac{n_{1}(B)\,n_{1}(A\backslash B)}{n_{1}(A)\,n_{0}(B)^{2}}}$
where $A\backslash B$ denotes the complement of subregion $B$. See Section 7.1.3 for details.
4. Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is consistent. A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous Poisson processes. This assumption is necessary to derive the standard errors.
In this section, we relax both assumptions. Since neither can be relaxed from the data alone, we consider additional information. In Section 4.1, we examine a generalization of the ratio estimator in which sampling weights adjust for differential nonresponse. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets.
4.1 Representativeness
The main assumption in Section 3 is that the status and spatial distribution of the respondents is representative. If this assumption holds, then the ratio estimator is consistent for a wider class of point processes in which Campbell’s Theorem applies, not just the Poisson process. See Daley and Vere-Jones (2007) for a general statement of Campbell’s Theorem. If this assumption is violated, then $\hat{\Lambda}_{0}(B)$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(B)\,/\,n_{1}(B)$ would not be representative, and $\hat{\Lambda}_{0}(B)$ would likely underestimate $\Lambda_{0}(B)$, even in large samples.
We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid as discussed in Section 2. Nevertheless, to evaluate how the estimates might change when this assumption is violated, we generalize the estimator in Section 3.2 to accommodate spatially varying sampling weights. The generalization is useful for determining the sensitivity of our estimates to differential nonresponse as we demonstrate in Section 5.1.
Let $w_{0}(x)$ and $w_{1}(x)$ denote two spatially varying weight functions. We define the weighted counts
$n^{w}_{0}(B)=\sum_{x\,\in\,\pi_{0}\cap B}w_{0}(x)\qquad\text{and}\qquad n^{w}_{1}(B)=\sum_{x\,\in\,\pi_{1}\cap B}w_{1}(x)$
where $\pi_{i}\cap B$ refers to the points of the process $\pi_{i}$ that fall in subregion $B\subseteq A$. We also define the weighted estimators for $\Lambda_{0}(A)$ and $p$,
$\hat{\Lambda}_{0}^{w}(A)=\dfrac{N_{1}(A)\,n^{w}_{0}(A)}{n^{w}_{1}(A)}\qquad\text{and}\qquad\hat{p}^{w}=\frac{n^{w}_{1}(A)}{N_{1}(A)}$
Under Campbell’s Theorem, the means $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)|N_{1}(A)]$ are
[…]
We assume the weight functions $w_{0}$ and $w_{1}$ can be chosen in such a way that the expected value is inversely proportional with the probability of a response. That is,
$p=\mathbb{E}[w_{0}(x)]\,p_{0}(x)=\mathbb{E}[w_{1}(x)]\,p_{1}(x)$
It follows that $\mathbb{E}[n^{w}_{0}(A)]$ and $\mathbb{E}[n^{w}_{1}(A)\,|\,N_{1}(A)]$ simplify to
$\mathbb{E}\left[n^{w}_{0}(A)\right]=p\,\Lambda_{0}(A)\qquad\text{and}\qquad\mathbb{E}\left[n^{w}_{1}(A_{i})\,|\,N_{1}(A)\right]=p\,N_{1}(A)$
and that $\hat{\Lambda}_{0}^{w}(A)$ is approximately normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}^{w}(A)]=\dfrac{N_{1}(A)\,n_{0}^{w}(A)}{n_{1}^{w}(A)}\sqrt{\dfrac{\hat{\mu}^{2}_{0}(A)+\hat{\sigma}_{00}(A,A)}{n_{0}^{w}(A)^{2}}+\dfrac{\hat{\mu}_{1}^{2}(A)}{n^{w}_{1}(A)^{2}}+\dfrac{(N_{1}(A)-1)\,\hat{\sigma}_{11}(A,A)}{N_{1}(A)\,n^{w}_{1}(A)^{2}}-\dfrac{1}{N_{1}(A)}-\dfrac{2\,\hat{\sigma}_{01}(A,A)}{n_{0}^{w}(A)\,n_{1}^{w}(A)}}$
where
$\hat{\mu}_{k}^{2}(A)=\sum_{x\,\in\,\pi_{k}\cap A}\mathbb{E}\bigl{[}w_{k}^{2}(x)\bigr{]}\qquad\text{and}\qquad\hat{\sigma}_{kl}(A,A)=\sum_{\begin{subarray}{c}x\,\in\,\pi_{k}\cap A\\[1.39998pt] y\,\in\,\pi_{l}\cap A\\[1.39998pt] x\neq y\end{subarray}}\mathbb{C}\bigl{(}w_{k}(x),\,w_{l}(y)\bigr{)}$
See Section 7.1.4 for details including derivations for the subregion and subtotal estimators.
The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error $\hat{\Lambda}_{0}(A)$ from Section 3.2 if $w_{0}(x)=w_{1}(x)=1$. In this case, $\hat{\mu}_{0}^{2}(A)=n_{0}(A)=n_{0}^{w}(A)$, $\hat{\mu}_{1}^{2}(A)=n_{1}(A)=n^{w}_{1}(A)$, and $\hat{\sigma}_{00}=\hat{\sigma}_{01}=\hat{\sigma}_{11}=0$.
Note that $\hat{\Lambda}_{0}(A)$ and $\hat{\Lambda}_{0}^{w}(A)$ are related by the relationship
$\hat{\Lambda}_{0}^{w}(A)=\frac{N_{1}(A)\ n^{w}_{0}(A)}{n^{w}_{1}(A)}=\hat{\Lambda}_{0}(A)\,\frac{n^{w}_{0}(A)\,/\,n_{0}(A)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
It follows that the relative bias depends only on the ratio of the average weights $n^{w}_{0}(A)\,/\,n_{0}(A)$ and $n^{w}_{1}(A)\,/\,n_{1}(A)$. For example, if the average permitted vendor is twice as likely to respond as the average unpermitted vendors, $n^{w}_{0}(A)\,/\,n_{0}(A)$ is twice the size of $n^{w}_{1}(A)\,/\,n_{1}(A)$, and $\hat{\Lambda}_{0}(A)$ is twice the size of $\hat{\Lambda}_{0}^{w}(A)$.
The estimators $\hat{\tau}(A)$ and $\hat{\tau}^{w}(A)$ are similarly related by the relationship
$\hat{\tau}^{w}(A)=\frac{N_{1}(A)\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)}{n^{w}_{1}(A)}=\hat{\tau}(A)\,\frac{\left(n^{w}_{0}(A)+n^{w}_{1}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)}{n^{w}_{1}(A)\,/\,n_{1}(A)}$
The same is true for subregion and subtotal estimators, see Section 7.1.4.
4.2 Overdispersion
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because $\hat{\Lambda}_{0}(A)$ is consistent under the wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated.
Suppose vendors serve customers who congregate at fixed locations, which we call markets. Let $M(B)$ denote the number of markets in any subregion $B\subseteq A\subset\mathbb{R}^{2}$. If every market is in a competitive equilibrium such that the expected profit for a new vendor is the same at each market, then the time between vendor arrivals at each market may be well-approximated by an exponential distribution with a constant arrival rate $\theta_{i}$. Further suppose each market starts with one vendor at $t=0$. Then at the time the survey was conducted ($t=1$), the number of vendors, $N_{i}(B)$, would follow a Poisson distribution with mean $\Lambda_{i}(B)=\theta_{i}\,M(B)$, and the results of Section 3.2 hold.
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process. That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\theta_{i}\,N^{t}_{i}(B)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\Lambda_{i}(B)=\exp(\theta_{i})\,M(B)$ and variance $\Lambda_{i}(B)\bigr{(}\Lambda_{i}(B)-M(B)\bigr{)}\,/\,M(B)$. i.e.,
$N_{i}(B)\sim\ \text{Negative Binomial}\left(\Lambda_{i}(B),\ \Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
where we parameterize the negative binomial distribution by its mean and variance for comparison with Section 3.2. In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details.
To derive the standard errors, suppose a portion of the markets, $p$, were selected for the survey at random so that
$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\Lambda_{i}(B),\ p\,\Lambda_{i}(B)\,\frac{\Lambda_{i}(B)-M(B)}{M(B)}\right)$
The conditional distribution of $n_{1}(B)$ given $N_{1}(A)$ is then
$n_{1}(B)\,|\,N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,q(B)\,M(A)\bigl{)}$
The conditional mean and variance of $n_{1}(A)$ given $N_{1}(A)$ is
$\mathbb{E}\left[n_{1}(A)\,|\,N_{1}(A)\right]=p\,N_{1}(A)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(A)\,|\,N_{1}(A)\bigr{)}=u_{1}(A)\,p\,(1-p)\,N_{1}(A)$
so that under increasing domain asymptotics, the ratio estimator
$\hat{\Lambda}_{0}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
is asymptotically normal with mean $\Lambda_{0}(A)$ and standard error
$\hat{\text{SE}}[\hat{\Lambda}_{0}(A)]=\dfrac{N_{1}(A)\,n_{0}(A)}{n_{1}(A)}\sqrt{\hat{u}_{0}(A)\left(\dfrac{1}{n_{0}(A)}\right)+u_{1}(A)\left(\dfrac{1}{n_{1}(A)}-\dfrac{1}{N_{1}(A)}\right)}$
where
$\hat{u}_{0}(B)=\cfrac{\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}-M(B)}{M(B)}\qquad\text{and}\qquad u_{1}(A)=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.1.5 for details including derivations for the subregion and subtotal estimators. Note that our parameterization makes use of the relationship $q(B)=M(B)\,/\,M(A)$.
The standard error derived in this section for $\hat{\Lambda}_{0}(A)$ is identical to the standard error for $\hat{\Lambda}_{0}(A)$ from Section 3.2 except for the weights $u_{0}$ and $u_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market. For example, if there are five vendors in each market on average, then $u_{0}$ and $u_{1}$ are approximately 4 and the standard error of the ratio estimator derived in this section is twice the size of the standard error derived in Section 3.2. The same is true for subregion and subtotal estimators, see Section 7.1.5.
5. Data Analysis
We apply the results presented in Section 3 and 4 to the data described in Section 2. In Section 5.1, we use the ratio estimator described in Section 3.1 and 3.2 to estimate the number of street vendors in New York City. In Section 5.2, we use the extensions described in Section 4.1 and 4.2 to deteremine the sensitivity of the estimate to the underlying assumptions. Sensitivity is assessed using additional data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03) and the American Community Survey from the U.S. Census Bureau (accessed 2022-12-22).
5.1 Estimation
@@
According to the survey data described in Section 2.2, $n_{1}(A)=349$ mobile food vendor respondents indicated they had a permit, and $n_{0}(A)=1,051$ indicated they did not. Since there are $N_{1}(A)=5,100$ permits available, it follows that the number of vendors without permits is 15,3505,100×1,051/ 349formulae-sequence153505100105134915,350$\approx$ 5,100$\times$ 1,051\,/\,34915 , 350 ≈ 5 , 100 × 1 , 051 / 349. Combined with the $5,100$ permitted vendors, the total number of mobile food vendors is approximately $20,500$.
@@
Estimates for select subregions are listed in Table 1. We find that approximately one quarter of street vendors are located in the Manhattan neighborhoods of Chelsea, Clinton, and Lower Manhattan. Another quarter is located in West and North Queens. We also report the margin of error (i.e., two standard errors, half of the width of a 95% confidence interval).
@@
These estimates can be checked by comparing them to a second, independent assessment conducted by the Street Vendor Project in the Bronx on May 13 and 15, 2022, in which the number of street vendors was documented through on-the-ground observations. The independent assessment identified 188 total street vendors near Fordham Road (Fordham Road BID and Street Vendor Project 2024). According to the survey data, 17 respondents indicated that they operate near Fordham Road, yielding an estimate of 166 street vendors total with a standard error of 54. We conclude this estimate is consistent with the 188 street vendors counted independently by Fordham Road BID and Street Vendor Project (2024).
{ThreePartTable}{TableNotes}
This table provides estimates for the number of mobile food vendors and non-veteran general merchandise vendors. The number of veteran general merchandise vendors is approximately 1,000. First amendment vendors and vendors in NYC Parks are not included.
Table 1: Number of Street Vendors
Respondents Population Margin of Error
Bronx
   Bronx Park and Fordham 68 757 239
   Southeast Bronx 28 385 160
   High Bridge and Morrisania 20 233 127
   Central Bronx 9 96 81
   Other Bronx 31 264 137
   Total 156 1,735 376
Brooklyn
      Northwest Brooklyn 46 613 204
      Bushwick and Williamsburg 41 433 175
      Sunset Park 26 344 153
                  Flatbush 19 242 128
            Borough Park 10 134 94
            Southwest Brooklyn 6 64 66
            Other Brooklyn 28 374 159
            Total 176 2,205 418
Manhattan
            Chelsea and Clinton 341 3,289 490
         Lower Manhattan 105 1,250 287
         Gramercy Park and Murray Hill 97 1,062 274
         Lower East Side 65 808 232
         Greenwich Village and Soho 61 726 222
         Upper East Side 56 534 194
         Central Harlem 46 483 185
         Inwood and Washington Heights 33 459 176
         Upper West Side 48 429 172
         East Harlem 38 425 174
         Total 890 9,464 879
Queens
         West Queens 396 4,650 692
      North Queens 67 896 251
      Northwest Queens 23 312 144
      Jamaica 14 181 110
      Other Queens 31 406 168
      Total 531 6,445 860
Staten Island
      Total 4 58 58
New York City
   Total 1,905 21,857 1,941
@@
Figure 1: Location of Street Vendors
[Uncaptioned image][Uncaptioned image]
[Uncaptioned image][Uncaptioned image]
These maps compare the spatial distribution of respondents (left) to the distribution of vendors suggested by administrative data (New York City Office of Administrative Trials and Hearings (OATH), top right) and federal data (American Community Survey (ACS), bottom right).
5.2 Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. We believe trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
Nevertheless, it is possible that canvassing operations systematically missed locations. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S. Census Bureau (accessed 2022-12-22). These comparisons are visualized in Figure 1. We find the spatial distribution of vending locations largely agree (top panels) as do the spatial distribution of the residencies (bottom panels). This suggests that, if there were a region of street vendors missed by the SVP, it was also missed by law enforcement and the ACS.
We can also combine the data from OATH and ACS with the results from Section 4.1 to assess the sensitivity of the estimates in Section 5.1 to differential nonresponse, as we now demonstrate. While the difference between these new estimates and the estimates provided in Section 5.1 are not statistically significant at the 95% level, we believe they provide insight into how the estimates in Section 5.1 might theoretically err.
First, we consider the scenario in which the probability a vendor responds is inversely proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code each day, as reported in the OATH data for the year 2021. This might be the case if vendors in high enforcement areas are reluctant to interact with survey takers. The weights are the inverse probabilities—the average number of enforcement actions taken against vendors in the surrounding ZIP Code. We find the weighted counts for food vendors are $n^{w}_{0}(A)=25.7$ and $n^{w}_{1}(A)=5.24$. This suggests there are $\left(n_{0}^{w}(A)+n_{1}^{w}(A)\right)\,/\,\left(n_{0}(A)+n_{1}(A)\right)\,/\left(n_{1}^{w}(A)\,/\,n_{1}(A)\right)=47\%$ more food vendors than estimated in Section 5.1. A similar argument for merchandise vendors suggests there are 9% more merchandise vendors than estimated in Section 5.1. Put together, the weights suggest the number of street vendors in New York City is 40% higher than estimated in Section 5.2.
The second scenario we consider is one in which the probability a vendor responds in proportional to the average number of vendors cited for violating street vending laws in the surrounding ZIP Code. This might be the case if vendors in high enforcement areas are easier to identify and thus easier to survey. Under this scenario, there are 8% fewer food vendors than estimated in Section 5.1 and 9% fewer merchandise vendors, suggesting the number of street vendors in New York City could be 8% lower than estimated in Section 5.2.
The final two scenarios are identical to the first two scenarios, except we use the number of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in the surrounding Public Use Microdata Area as estimated from the 2017-2021 U.S. Census Bureau (accessed 2022-12-22). The result is less extreme than using the average number of vendors cited. If the probability of response is inversely proportional to this estimate, there are 33% more vendors New York City than estimated in Section 5.2. If the probability of response is proportional to this estimate, there are 5% fewer vendors. Combining these four scenarios suggests a conservative range of between 21,000 and 32,200 street vendors operate in New York City.
The second assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. We argue that this assumption is reasonable because a wide class of data generating processes are well-approximated by the Poisson process under the law of rare events as discussed in Section 3.2.